# A universal upper bound on the photon sphere radius in higher-dimensional black holes

### v1

@@ Abstract
In this work, we derive a universal upper bound for the photon sphere radius in static, spherically symmetric, asymptotically flat black hole spacetimes of arbitrary dimension $n\geq 4$, in the presence of an anisotropic matter field satisfying the weak energy condition and a non-positive trace of the energy-momentum tensor. Using the effective potential method, we obtain the bound $r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}$, where $M$ is the ADM mass of the black hole. This bound reduces to $r_{\gamma}\leq 3M$ in four dimensions, consistent with the known result in the literature. Our result provides a dimension-dependent upper bound for photon spheres and deepens the understanding of spacetime structure in higher-dimensional gravitational theories.
@@ I Introduction
For the four-dimensional Schwarzschild black hole, the photon sphere is located at $r_{\gamma}=3M$. For more general, static, spherically symmetric, and asymptotically flat black holes surrounded by matter fields (hairy black holes), it was proven by Hod that the photon sphere radius satisfies the universal bound $r_{\gamma}\leq 3M$, provided the matter obeys the weak energy condition and its energy-momentum tensor has a non-positive trace Hod:2013jhd . This bound is saturated by the vacuum (bald) Schwarzschild black hole. Subsequent studies have explored related upper and lower bounds for photon spheres of compact stars Hod:2020pim ; Peng:2020shc ; Liu:2024odv ; Hod:2023jmx ; Song:2025fgq .
Given the significant interest in higher-dimensional theories of gravity, such as string theory and brane-world models, a natural and important question arises: does a similar universal bound exist for the photon sphere in higher-dimensional black hole spacetimes? Understanding how spacetime dimensionality affects fundamental structures like the photon sphere is essential for probing the geometry of higher-dimensional gravity and identifying potential observational signatures of extra dimensions.
In this work, we extend Hod’s analysis to arbitrary dimensions $n\geq 4$. Within a model-independent framework for static, spherically symmetric, asymptotically flat black holes with anisotropic matter, we derive the characteristic equation for the photon sphere. Under the assumptions of the weak energy condition and a non-positive trace for the energy-momentum tensor, we prove the existence of a photon sphere and establish the universal upper bound:
@@
In four dimensions ($n=4$), this reduces to $r_{\gamma}\leq 3M$, confirming Hod’s result Hod:2013jhd . Our bound is dimension-dependent and suggests that the presence of matter fields tends to pull the photon sphere inward compared to the corresponding vacuum (Tangherlini) black hole. This work generalizes a key four-dimensional result to higher dimensions, deepening our understanding of black hole structure in extended theories of gravity.
The paper is organized as follows. In Sec. II, we describe the general higher-dimensional black hole spacetime and the corresponding Einstein field equations for an anisotropic fluid. Sec. III contains our main analysis: we derive the photon sphere condition, prove its existence, and establish the universal upper bound under appropriate energy conditions. We conclude with a summary and discussion of implications in Sec. IV. Throughout, we use natural units with $G=c=1$.
@@ II DESCRIPTION OF THE SYSTEM IN THE HIGHER-DIMENSIONAL BLACK HOLE
In this work, we focus on the static, spherically symmetric, and asymptotically flat higher-dimensional black hole systems. The $n$-dimensional Einstein-Hilbert action with matter fields is given by
@@
Taking $T^{t}{}_{t}=-\rho$$T^{r}{}_{r}=p_{r}$ and $T^{\theta}{}_{\theta}=T^{\phi}{}_{\phi}=p_{t}$ , where $\rho$, $p$ and $p_{T}$ are identified as the energy density, radial pressure, and tangential pressure, respectively. The static diagonal energy-momentum tensor of an anisotropic fluid can be expressed as
$T^{\mu}{}_{\nu}=\mathrm{diag}(-\rho(r),\,p_{r}(r),\,\underbrace{p_{t}(r),\,p_{t}(r),\,\ldots,\,p_{t}(r)}_{n-2\ \text{terms}})\;.$ (19)
From the Einstein field equations $G^{\mu}{}_{\nu}=8\pi T^{\mu}{}_{\nu}$ and using Eq. (11), we obtain
@@
$\delta^{\prime}=-\frac{8\pi r(\rho+p_{r})}{(n-2)\mu}\;.$ (21)
@@
The condition of finite mass configuration characteristics implies
@@
$p_{r}^{\prime}=\frac{2T+(n-1)\rho-(n+1)p_{r}}{2r}-(\frac{n-3}{2r}+\frac{8\pi rp_{r}}{n-2})\frac{\rho+p_{r}}{\mu}\;,$ (27)
@@
is the trace of the energy momentum tensor $T^{\mu}{}_{\nu}$. Eq. (27) reduces to the four-dimensional case in Ref. Hod:2013jhd .
III Upper bound on the radius of the photon sphere of black holes
@@
It can be directly noted that this reduces to the four-dimensional case Hod:2013jhd .
@@
    Weak Energy Condition (WEC): The components of the energy-momentum tensor satisfy the weak energy condition (WEC). This means that the energy density of the matter fields is positive semidefinite, i.e.,
    $\rho\geq 0\;,$ (44)
    and that it bounds the pressures, which implies the inequality
    $\rho+p_{r}\geq 0\;,\quad\mathrm{and}\quad\rho+p_{t}\geq 0\;.$ (45)
@@
Next, we analyze the behavior of $P(r)$ in the region from the event horizon to the photon sphere. By substituting the photon sphere characteristic inequality (42), together with the energy conditions (45) and (28) for the matter field, into the pressure gradient (43), we obtain
@@
for the photon sphere. Since $m(r)$ is a non-decreasing function of $r$ and $m(r_{\gamma})\leq M$, where $M=m(r\to\infty)$ is the total ADM mass of the black hole spacetime, we finally obtain the universal upper bound
@@
It is worth emphasizing that, for static, spherically symmetric and asymptotically flat spacetimes of arbitrary dimension ($n\geq 4$) which satisfy the weak energy condition (45) and the non-positive trace condition (46), a universal upper bound for the photon sphere can be directly obtained from Eq. (54), without the need for further complicated calculations. Moreover, since the boundary of a black hole shadow is determined by the photon sphere, this result can also provide a theoretical benchmark for the observation of black hole shadows, potentially constraining extra dimensions.
IV Discussion and conclusion
In the context of general relativity and higher-dimensional gravitational theories, the photon sphere of a black hole, as a key structure influencing its shadow and dynamics, holds significant importance for establishing universal bounds on its radius. Although the relevant bounds in four-dimensional spacetime have been thoroughly discussed, it remains unclear whether similar universal upper bounds exist for the photon spheres of higher-dimensional black holes.
In this work, we perform an analysis of the photon spheres for static, spherically symmetric and asymptotically flat higher-dimensional black holes. In particular, under the constraints of the weak energy condition and the non-positive trace of the energy-momentum tensor, we obtain the upper bound $r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}$ for the radius of its photon sphere via an analytical method, where $n$ is the spacetime dimension and $M$ is the total ADM mass of the spacetime. It is worth noting that we have not restricted the specific form of $\mu(r)$ in the derivation process, therefore, this is a dimension-dependent universal upper bound. Moreover, the upper bound of the photon sphere radius for an arbitrary spacetime dimension can be directly obtained from Eq. (54), which simplifies the calculations involved in the conventional effective potential method.
Our result generalizes a fundamental geometric constraint from four to higher dimensions, strengthening the theoretical toolkit for analyzing black holes in theories beyond standard general relativity, such as string theory and brane-world models. The bound suggests that observational measurements of black hole shadow sizes, for instance by next-generation very-long-baseline interferometry, could in principle be used to test for signatures of extra dimensions by comparing the inferred photon sphere radius with the dimension-dependent upper limit.
Future research could extend this work in several promising directions. A natural and challenging extension would be to study rotating (Kerr-like) higher-dimensional black holes, investigating whether a similar bound exists and how it depends on the spin parameter. Exploring photon sphere bounds in asymptotically de Sitter or anti-de Sitter spacetimes would also be valuable, especially for applications in the gauge/gravity duality and cosmological contexts. Furthermore, understanding the precise saturation condition of the bound in higher dimensions and its relation to specific matter models or no-hair theorems presents an interesting theoretical challenge. Finally, further investigation into the relationship between the photon sphere bound and other black hole characteristics, such as quasinormal mode spectra, thermodynamic stability, or holographic complexity, could reveal deeper geometric principles.
In conclusion, we have derived a universal, dimension-dependent upper bound for the photon sphere radius in a broad class of higher-dimensional black holes. This work not only consolidates our understanding of black hole geometry across dimensions but also provides a new theoretical tool for probing the structure of spacetime in extended theories of gravity.

### v2

# Bounds on the photon sphere radius for spherically symmetric black holes in $\bm{n}$-dimensional Einstein gravity

@@ Abstract
The photon sphere, a hypersurface of circular null geodesics, plays a fundamental role in characterizing black hole spacetimes, influencing phenomena such as black hole shadows, gravitational lensing, and quasinormal modes. In this work, we derive both upper and lower bounds on the photon sphere radius for static, spherically symmetric, asymptotically flat black holes within $n$-dimensional Einstein gravity ($n\geq 4$), assuming an anisotropic matter field satisfying the weak energy condition and a non‑positive trace of the energy‑momentum tensor. For the upper bound we obtain $r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}$, where $M$ is the ADM mass, which reduces to $r_{\gamma}\leq 3M$ for $n=4$ in agreement with Hod’s result. For the lower bound, under the additional assumption that $|r^{n-1}p_{r}(r)|$ is monotonically decreasing, we prove $r_{\gamma}\geq(\frac{n-1}{2})^{1/(n-3)}r_{H}$, where $r_{H}$ is the radius of the outer event horizon; for $n=4$ this gives $r_{\gamma}\geq\frac{3}{2}r_{H}$, also consistent with Hod’s four‑dimensional theorem. These results provide dimension‑dependent geometric constraints that generalize well‑known four‑dimensional bounds to a specific class of higher‑dimensional black holes (described by a Tangherlini‑type metric) and deepen our understanding of spacetime structure in higher‑dimensional gravitational theories.
@@ I Introduction
For the four-dimensional Schwarzschild black hole, the photon sphere is located at $r_{\gamma}=3M$. For more general static, spherically symmetric, asymptotically flat black holes surrounded by matter, Hod proved that the photon sphere radius satisfies the upper bound $r_{\gamma}\leq 3M$ provided the matter obeys the weak energy condition and its energy‑momentum tensor has a non‑positive trace Hod:2013jhd . This bound is saturated by the vacuum (bald) Schwarzschild black hole. Subsequently, Hod also established a lower bound $r_{\gamma}\geq\frac{3}{2}r_{H}$ under the additional condition that $|r^{3}p_{r}(r)|$ decreases monotonically Hod:2020pim . Subsequent studies have explored related upper and lower bounds for photon spheres of compact stars Lu:2019zxb ; Peng:2020shc ; Liu:2024odv ; Hod:2023jmx ; Song:2025fgq . These bounds are not only of theoretical interest but also provide insights into the possible size of black hole shadows and the structure of hairy black holes.
Given the significant interest in higher-dimensional theories of gravity-motivated by string theory, brane‑world models, and the gauge/gravity duality-a natural question arises: do similar bounds exist for the photon sphere in higher‑dimensional black hole spacetimes? Understanding how spacetime dimensionality affects fundamental structures like the photon sphere is essential for probing the geometry of higher-dimensional gravity.
In this work, we fill this gap by deriving both an upper and a lower bound for the photon sphere radius in static, spherically symmetric, asymptotically flat black holes of $n$-dimensional Einstein gravity ($n\geq 4$). Our analysis is performed within a model‑independent framework that assumes a Tangherlini‑type metric structure and an anisotropic matter fluid satisfying the weak energy condition and a non‑positive trace of the energy‑momentum tensor. For the upper bound, we obtain
@@
which for $n=4$ reduces to $r_{\gamma}\leq 3M$. For the lower bound, we impose one further physically motivated condition-the monotonic decrease of $|r^{n-1}p_{r}(r)|$-and prove
$r_{\gamma}\geq\bigg(\frac{n-1}{2}\bigg)^{1/(n-3)}r_{H}\;,$ (2)
which for $n=4$ becomes $r_{\gamma}\geq\frac{3}{2}r_{H}$. These results provide a complete set of dimension‑dependent geometric constraints that generalise the well‑known four‑dimensional bounds to a specific class of higher‑dimensional black holes.
The paper is organized as follows. In Sec. II, we describe the general higher-dimensional black hole spacetime and the corresponding Einstein field equations for an anisotropic fluid. In Sec. III, we derive the photon sphere condition, prove its existence, and establish the upper bound under appropriate energy conditions. In Sec. IV, we derive the lower bound, employing the monotonicity property of the radial pressure function. We conclude with a discussion of the results, their limitations, and possible future extensions in Sec. V. In App. A, we provides a rigorous proof of an inequality needed in the lower bound analysis. Throughout, we use natural units with $G=c=1$.
@@ II DESCRIPTION OF THE SYSTEM IN THE HIGHER-DIMENSIONAL BLACK HOLE
In this work, we focus on static, spherically symmetric, and asymptotically flat higher-dimensional black hole systems within Einstein gravity. The $n$-dimensional Einstein-Hilbert action with matter fields is given by
@@
The matter field is modelled as an anisotropic fluid, whose energy‑momentum tensor is diagonal in the static orthonormal frame:
$T^{\mu}{}_{\nu}=\mathrm{diag}(-\rho(r),\,p_{r}(r),\,\underbrace{p_{t}(r),\,p_{t}(r),\,\ldots,\,p_{t}(r)}_{n-2\ \text{terms}})\;,$ (20)
where $\rho$ is the energy density, $p_{r}$ is the radial pressure, and $p_{t}$ is the tangential pressure. From the Einstein field equations $G^{\mu}{}_{\nu}=8\pi T^{\mu}{}_{\nu}$ and using Eq. (12), we obtain
@@
$\delta^{\prime}=-\frac{8\pi r(\rho+p_{r})}{(n-2)\mu}\;,$ (22)
@@
The condition for a finite mass configuration characteristics implies
@@
$p_{r}^{\prime}=\frac{2T+(n-1)\rho-(n+1)p_{r}}{2r}-\bigg(\frac{n-3}{2r}+\frac{8\pi rp_{r}}{n-2}\bigg)\frac{\rho+p_{r}}{\mu}\;,$ (28)
@@
is the trace of the energy momentum tensor $T^{\mu}{}_{\nu}$.
III Upper bound on the photon sphere radius
@@
    Weak Energy Condition (WEC): The energy density of the matter fields is positive semidefinite, and that it bounds the pressures
    $\rho\geq 0\quad\;;\quad\rho\geq|p_{r}|,|p_{T}|\;.$ (45)
@@
Next, we analyze the behavior of $P(r)$ in the region from the event horizon to the photon sphere. By substituting the photon sphere characteristic inequality (43), together with the energy conditions (LABEL:weak2) and (29) for the matter field, into the pressure gradient (44), we obtain
@@
for the photon sphere. Since $m(r)$ is a non-decreasing function of $r$ and $m(r_{\gamma})\leq M$, where $M=m(r\to\infty)$ is the total ADM mass of the black hole spacetime, we finally obtain the upper bound
@@
IV Lower bound on the photon sphere radius
To obtain the lower bound, we need an additional condition:
  • (3).
    Monotonicity of the radial pressure function: The radial pressure function
    $P\equiv|r^{n-1}p_{r}(r)|$ (55)
    decreases monotonically with $r$. This condition generalizes the four-dimensional assumption used by Hod Hod:2020pim and is motivated by the behavior of matter fields in many known hairy black hole solutions, where the radial pressure typically decays rapidly outside the horizon. In the vacuum region, $p_{r}=0$, so the condition is trivially satisfied.
Following the Ref. Hod:2020pim , we first derive a lower bound on the mass of the matter fields exterior to the horizon in higher-dimensional black holes. From Eq. (24) and condition (45), one obtains
$m_{\mathrm{ex}}=\frac{8\pi}{n-2}\int_{r_{H}}^{\infty}r^{n-2}\rho(r)dr\geq-\frac{8\pi}{n-2}\int_{r_{H}}^{\infty}r^{n-2}p_{r}(r)dr\;.$ (56)
Crucially, it has been shown that under the conditions (45) and (46), the function $P(r)\equiv r^{n}p_{r}$ is non-positive and monotonically decreasing in the interval $r\leq[r_{H},r_{\gamma}]$:
$\{p_{r}(r)\leq 0\quad\mathrm{and}\quad(r^{n}p_{r})^{\prime}\leq 0\}\quad\mathrm{for}\quad r_{H}\leq r\leq r_{\gamma}\;.$ (57)
where $r_{\gamma}$ is the radius of photon sphere. Consequently,
$0\leq-r^{n}_{H}p_{r}(r_{H})\leq-r^{n}p_{r}(r)\quad\mathrm{for}\quad r_{H}\leq r\leq r_{\gamma}\;.$ (58)
Using the relation (58) in Eq. (56) gives the chain of inequalities
$m_{\mathrm{ex}}\geq-\frac{8\pi}{n-2}\int_{r_{H}}^{r_{\gamma}}r^{n-2}p_{r}(r)dr\geq-\frac{8\pi}{n-2}\int_{r_{H}}^{r_{\gamma}}\frac{r_{H}^{n}p_{r}(r_{H})}{r^{2}}dr=-\frac{8\pi}{n-2}r^{n}_{H}p_{r}(r_{H})\bigg(\frac{1}{r_{H}}-\frac{1}{r_{\gamma}}\bigg)$ (59)
for the mass $m_{\mathrm{ex}}$ of the external matter fields.
The location of the photon sphere is determined by the condition (38)
$\mathcal{R}(r_{\gamma})=3-n+\mu(n-1)-\frac{16\pi r_{\gamma}^{2}p_{r}(r_{\gamma})}{n-2}=0\;.$ (60)
Substituting the lower bound (59) into Eq. (60) and using the relation (25), one obtains the inequality
$2-(n-1)\frac{r^{n-3}_{H}}{r^{n-3}_{\gamma}}-\frac{16\pi r^{2}_{\gamma}p_{r}(r_{\gamma})}{n-2}+\frac{16\pi(n-1)}{(n-2)r^{n-3}_{\gamma}}r^{n}_{H}p_{r}(r_{H})\bigg(\frac{1}{r_{H}}-\frac{1}{r_{\gamma}}\bigg)\geq 0\;,$ (61)
which characterizes the null circular geodesics of the spherically-symmetric black-hole spacetime. In addition, using the inequality $r^{n-1}_{H}p_{H}\leq r^{n-1}_{\gamma}p_{\gamma}$, which follows from the assumed monotonic behavior of the radial pressure function (55) and the fact that the radial pressure is non-positive between the black-hole horizon and the innermost null circular geodesic, one deduces from (61) the inequality
$2-(n-1)\frac{r^{n-3}_{H}}{r^{n-3}_{\gamma}}-\frac{16\pi r^{2}_{\gamma}p_{r}(r_{\gamma})}{n-2}+\frac{16\pi(n-1)}{(n-2)r^{n-3}_{\gamma}}r_{H}r_{\gamma}^{n-1}p_{r}(r_{\gamma})\bigg(\frac{1}{r_{H}}-\frac{1}{r_{\gamma}}\bigg)\geq 0\;,$ (62)
Simplifying the above equation, we have
$2-(n-1)\frac{r^{n-3}_{H}}{r^{n-3}_{\gamma}}+\frac{16\pi r^{2}_{\gamma}p_{r}(r_{\gamma})}{n-2}\bigg[n-2-(n-1)\frac{r_{H}}{r_{\gamma}}\bigg]\geq 0\;.$ (63)
Substituting the condition (60) in to Eq. (63), we have
$2-(n-1)\frac{r^{n-3}_{H}}{r^{n-3}_{\gamma}}+[(n-1)\mu(r_{\gamma})-(n-3)]\bigg[n-2-(n-1)\frac{r_{H}}{r_{\gamma}}\bigg]\geq 0\;.$ (64)
Introducing $x=\frac{r_{\gamma}}{r_{H}}>1$, so $r^{n-3}_{\gamma}/r^{n-3}_{H}=x^{n-3}$, and let $\mu_{\gamma}=\mu(r_{\gamma})$. Then Eq. (64) becomes
$2-(n-1)x^{-(n-3)}+[(n-1)\mu_{\gamma}-(n-3)]\bigg[n-2-(n-1)x^{-1}\bigg]\geq 0\;.$ (65)
Define
$G(x)=2-(n-1)x^{-(n-3)}+[(n-1)\mu_{\gamma}-(n-3)]\bigg[n-2-(n-1)x^{-1}\bigg]\;.$ (66)
So that $G(x)\geq 0$. Since $p_{r}(r_{\gamma})\leq 0$, Eq. (60) implies
$A\equiv(n-1)\mu_{\gamma}-(n-3)\leq 0\;.$ (67)
We now analyse the function $G(x)$ for $x>1$. Define
$f(x)=2-(n-1)x^{-(n-3)}\;,\quad g(x)\equiv n-2-(n-1)x^{-1}\;,$ (68)
so that $G(x)=f(x)+Ag(x)$ with $A\leq 0$ .
Consider two cases according to the sign of $g(x)$.
  • (1).
    $g(x)\geq 0$ (i.e., $x\geq\frac{n-1}{n-2}$): Then $Ag(x)\leq 0$ and consequently $G(x)\leq f(x)$. The condition $G(x)\geq 0$ therefore requires $f(x)\geq 0$, i.e.,
    $2-(n-1)x^{-(n-3)}\geq 0\;,$ (69)
    Hence, in this case we must have
    $x\geq\bigg(\frac{n-1}{2}\bigg)^{\frac{1}{n-3}}\;.$ (70)
  • (2).
    $g(x)<0$ (i.e. $1<x<\frac{n-1}{n-2}$): we prove that in this case, one actually always has $G(x)<0$. This is done by observing that
    $\frac{\partial G}{\partial A}=g(x)<0\;,$ (71)
    hence $G(x)$ attains its maximum at the smallest allowed value of $A$. The smallest possible $A$ is $A_{\mathrm{min}}=-(n-3)$ (corresponding to $\mu_{r}=0$). Substituting $A_{\mathrm{min}}=-(n-3)$ into Eq. (66) gives
    $G_{\mathrm{max}}(x)=(n-1)[-(n-4)-x^{-(n-3)}+(n-3)x^{-1}]\;.$ (72)
    For $n\geq 5$, $G_{\mathrm{max}}(x)$ is strictly negative for all $x>1$ (it vanishes at $x=1$ and is decreasing [^4] From $G_{\mathrm{max}}(x)=(n-1)[-(n-4)-x^{-(n-3)}+(n-3)x^{-1}]\;,$ (73) we have $G_{\mathrm{max}}(1)=0$, and $G^{\prime}_{\mathrm{max}}(x)=(n-1)[(n-3)x^{-(n-2)}-(n-3)x^{-2}]=(n-1)(n-3)x^{-2}\bigg(x^{-(n-4)}-1\bigg)\;.$ (74) For $n\geq 5$, we have $n-4\geq 1$ and $x^{-(n-4)}<1$ for $x>1$, so $G^{\prime}_{\mathrm{max}}(x)<0$.). For $n=4$, from Eq. (66) one obtains $G(x)=3\mu_{r}(2-3x^{-1})$, which is negative precisely when $1<x<\frac{3}{2}$. Thus, for any $n\geq 4$, if $g(x)<0$, then $G(x)<0$, which conflicts with Eq. (65).
Combining the two cases we conclude that from the condition $G(x)\geq 0$, we have
$x\geq\frac{n-1}{n-2}\;,\quad\mathrm{and}\quad x\geq\bigg(\frac{n-1}{2}\bigg)^{\frac{1}{n-3}}\;.$ (75)
Using the result (80) in appendix A, we obtain the necessary and sufficient condition for $G(x)\geq 0$:
$x\geq\bigg(\frac{n-1}{2}\bigg)^{\frac{1}{n-3}}\;.$ (76)
Returning to the original variables, this inequality becomes
$r_{\gamma}\geq\bigg(\frac{n-1}{2}\bigg)^{\frac{1}{n-3}}r_{H}\;,$ (77)
which provides a lower bound on the photon-sphere radius in higher-dimensional black holes.
V Discussion and conclusion
In this work, we derived both an upper and a lower bound for the photon sphere radius of static, spherically symmetric, asymptotically flat black holes in $n$-dimensional Einstein gravity ($n\geq 4$), under physically motivated assumptions on the matter fields. The upper bound,
$r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}\;,$ (78)
generalises Hod’s four‑dimensional result Hod:2013jhd and is saturated by the vacuum Tangherlini black hole. The lower bound,
$r_{\gamma}\geq(\frac{n-1}{2})^{1/(n-3)}r_{H}$ (79)
requires the additional condition that $|r^{n-1}p_{r}(r)|$ decreases monotonically; it reduces to $r_{\gamma}\geq\frac{3}{2}r_{H}$ for $n=4$, in agreement with Hod:2020pim . Together, these bounds provide a complete characterisation of the possible location of the photon sphere in this class of spacetimes.
It is important to emphasize the scope and limitations of our result. Our derivation relies on the Tangherlini-type metric ansatz (16), which assumes a direct product structure with a maximally symmetric transverse space and strict asymptotic flatness in all $n$ dimensions. This framework, while natural for a first generalization, excludes more intricate geometries such as warped spacetimes, brane-world scenarios, and models with compactified extra dimensions. In such settings, the properties of the photon sphere may differ substantially, and our bound may not apply.
Furthermore, we must address the potential observational implications-or lack thereof-for astrophysical black holes. In phenomenologically viable extra-dimensional models, the additional spatial dimensions are typically compactified at a scale far below the micrometer level, as constrained by current laboratory and astrophysical experiments. For solar-mass or supermassive black holes, whose horizons are of the order of kilometers to astronomical units, the effective gravitational dynamics at macroscopic scales reduce to four dimensions. Consequently, any corrections to the photon sphere radius arising from the presence of such small extra dimensions would be overwhelmingly negligible. Therefore, our result does not translate to observable deviations from the four-dimensional bound for astrophysical black holes, unless one entertains scenarios with non-compact or extremely large extra dimensions, which are already strongly disfavored by existing constraints.
Notwithstanding these limitations, our work provides a valuable theoretical extension within a well-defined class of higher-dimensional black holes. The bound offers a geometric constraint that may inform the study of black hole shadows in idealized higher-dimensional contexts and contributes to the toolkit for analyzing black holes in theories beyond standard four-dimensional general relativity.
Future research could extend this work in several promising directions. Possible extensions include investigating axisymmetric (rotating) configurations Cunha:2017qtt ; Cunha:2020azh , considering asymptotically de Sitter or anti-de Sitter spacetimes, or examining the implications of quantum corrections on the energy conditions. Such generalizations would further illuminate the interplay between energy conditions, horizon geometry, and the behavior of null geodesics in curved spacetimes.
Appendix A Proof of the inequality $\frac{n-1}{n-2}\leq(\frac{n-1}{2})^{1/(n-3)}$ for $n\geq 4$
We prove that for $n\geq 4$,
$\frac{n-1}{n-2}\leq\bigg(\frac{n-1}{2}\bigg)^{\frac{1}{n-3}}$ (80)
with equality if and only if $n=4$.
Set $x=n-1(\geq 3)$. Then the inequality becomes
$\frac{x}{x-1}\leq\bigg(\frac{x}{2}\bigg)^{\frac{1}{x-2}}\;.$ (81)
Taking natural logarithms on both sides, it is equivalent to
$\ln\bigg(\frac{x}{x-1}\bigg)\leq\frac{1}{x-2}\ln\bigg(\frac{x}{2}\bigg)\;,$ (82)
or, rearranging,
$(x-2)\ln\bigg(\frac{x}{x-1}\bigg)-\ln\bigg(\frac{x}{2}\bigg)\leq 0\;.$ (83)
Define
$F(x)\equiv(x-2)\ln\bigg(\frac{x}{x-1}\bigg)-\ln\bigg(\frac{x}{2}\bigg),\quad x\geq 3\;.$ (84)
we have $F(3)=0$. Its derivative is
$F^{\prime}(x)=\ln\bigg(\frac{x}{x-1}\bigg)-\frac{2x-3}{x(x-1)}\;.$ (85)
Using the elementary inequality $\ln(1+t)<t$ for $t>0$ with $t=\frac{1}{x-1}$, we obtain
$\ln\bigg(\frac{x}{x-1}\bigg)=\ln\bigg(1+\frac{1}{x-1}\bigg)<\frac{1}{x-1}\;.$ (86)
For $x>3$, it is easy to verify that
$\frac{1}{x-1}\leq\frac{2x-3}{x(x-1)}\;.$ (87)
Hence, for all $x>3$,
$\ln\bigg(\frac{x}{x-1}\bigg)<\frac{1}{x-1}<\frac{2x-3}{x(x-1)}\;,$ (88)
which implies $F^{\prime}(x)<0$ for $x>3$. Consequently, $F(x)$ is strictly decreasing on $3,\infty)$ and therefore $F(x)<F(3)=0$ for every $x>3$. This proves (80) with strict inequality for $x>3$ (i.e., for $n>4$), while equality holds at $x=3$ (i.e., at $n=4$).